Jim invests $30,000 in two separate accounts. One pays 4% each year simple interest the other pays 6% his investment earns 4620 over the three year period how much did he deposit in each account
Answer by jorel1380(3719)

Let n be the amount invested at 6%. Then the amount invested at 4% would be 30000-n. So:
3*.06n+3*.04(30000-n)=4620
.06n+.04(3000-n)=1540
.02n+1200=1540
.02n=340
n=17000
$17000 was invested at 6%; and $13000 was invested at 4%. ☺☺☺☺